OpenStudy (anonymous):

HELP PLEASE.. You have 2 m3 of a mineral ore sample. It has a mass of 1 × 104 kg. Calculate the density of the ore sample and express your results in g/cm3 with appropriate scientific notation.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

1*10^4 kg is equal to how many grams?

OpenStudy (anonymous):

10,000,000 grams.

jimthompson5910 (jim_thompson5910):

which is equivalent to 1*10^7 grams

jimthompson5910 (jim_thompson5910):

now convert 2 m^3 to cm^3

OpenStudy (anonymous):

How do I do that?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

1 m = 100 cm (1 m)^3 = (100 cm)^3 1 m^3 = 1,000,000 cm^3 1 m^3 = 1*10^6 cm^3

jimthompson5910 (jim_thompson5910):